OpenStudy (helpblahblahblah):

ILL GIVE MEDALS: 75 is what percent of 25?

OpenStudy (anonymous):

33 1/3%

OpenStudy (helpblahblahblah):

explain please

OpenStudy (ksaimouli):

x/100 = 25/75

OpenStudy (helpblahblahblah):

???

OpenStudy (ksaimouli):

x = 100(25/75)

OpenStudy (dulljackel09):

Its 18.75

OpenStudy (ksaimouli):

x = 100(1/3)

OpenStudy (ksaimouli):

x = 33 1/3%

OpenStudy (anonymous):

Ksaimouli explained it (: